# MobCLIP: Learning General-purpose Geospatial Representation at Scale

### v3

@@ 4.1.1 Multimodal Pretraining
We construct a store-level network based on consumer mobility from WeChat Pay, a nationwide electronic transaction platform in China operated by Tencent, with extensive offline payment coverage. Each offline store is mapped to an H3 grid cell based on its geographic coordinates. If a user makes purchases at stores in different grids within the same week, an undirected, weighted edge is formed between the corresponding grid cells, generating a weekly mobility graph. Aggregating data over 54 weeks and 50 million unique stores yields a large-scale graph at H3 resolution level 6, with approximately 200,000 nodes and 1.2 billion edges. This nationwide graph captures fine-grained human movement patterns and serves as the foundational mobility graph.
@@
More than 100 million Points-of-Interest (POIs) are collected from Tencent Maps, including both their names and category labels. POI names are typically short, referring to store names, company names, or types of public facilities. The POIs span 445 distinct categories[35]. For each POI, we concatenate its name and category, and encode the combined text using a pretrained text embedding model. In terms of images, we obtain Google Satellite imagery with a resolution of approximately 10 meters (at zoom level 14) for each H3 grid. Grid-level demographic statistics are processed by aggregating population structure data from WorldPop [44]. Each H3 grid records the population count by age group (5-year intervals) and gender in 2020.
@@
We split the dataset into 90% for training and 10% for validation to mitigate overfitting. The model is trained for 100 epochs with a batch size of 20,480, using the Adam optimizer with a learning rate of 0.001 and weight decay of 0.001 on a single NVIDIA A100 GPU for one hour. During training, we conduct a grid search over commonly used hyperparameters: learning rate $\in\{1e-4,2e-4,5e-4,1e-3,2e-3,5e-3\}$ and weight decay $\in\{1e-3,1e-2,1e-1\}$. A comprehensive list of tuned hyperparameters is provided in our supplementary GitHub repository. The LINE embedding for nationwide mobility graph was processed at Tencent’s Angel distributed graph computing platform [15], utilizing hundreds of computing cores and the embedding of over a billion edges is completed efficiently within a few hours.
4.1.2 General-purpose Evaluation
@@
To assess the generalizability of our model, we curated a set of 11 downstream tasks as shown in Table 1. These tasks were selected for their diversity and broad representativeness, covering three key domains: social, economic, and natural. They are evenly distributed across China, ensuring that the evaluation is free from geographic bias. In addition, considering the influence of statistical unit delineation in geospatial analysis, the downstream tasks are designed to span four distinct spatial scales — point, grid, county/district, and city/prefecture — to examine the impact of resolution effect in region representation learning. MobCLIP is trained at the H3 grid level. For point-level tasks, the embedding corresponding to the grid containing the point is extracted for evaluation. For administrative-level tasks, we compute the mean embedding of all grids within the administrative unit, which is then used to assess regression performance. This multi-scale evaluation further validates the generalizability of the model across varying spatial resolutions. The data sources are listed in the table and are primarily derived from WorldPop [4], MOSAIKS [32], and 2020 China Census Year Book [30]. Additionally, ENE (energy consumption) was collected from various local statistical yearbooks. The total energy consumption of one city/prefecture is calculated by converting different forms of energy usage into tonnes of standard coal equivalent. As this work has not yet been formally published, the data are not publicly available now. Furthermore, another unavailable dataset used in our evaluation is COS (offline consumption amount) from Tencent WeChat Pay, which was used to test whether the embeddings can predict regional consumption indices.
@@
Abbreviation Description Data Source Spatial Scale Publicly Available
@@
POP Population density WorldPop [4] Grid Y
EDU Average education years Census Year Book [30] County/District Y
ELD Elderly (65+) population ratio Census Year Book [30] County/District Y
HSR Hukou separation ratio Census Year Book [30] County/District Y
CRI Crime cases [55] Grid Y
@@
NTL Nighttime light intensity MOSAIKS [32] Point Y
HOU Per capita housing area Census Year Book [30] County/District Y
ENE Energy consumption City statistical yearbooks City/Prefecture N
COS Offline consumption amount Tencent Grid N
Natural Tasks
ELE Elevation MOSAIKS [32] Point Y
FOR Forest coverage MOSAIKS [32] Point Y
@@
We implement LightGBM[17] as the downstream prediction model. The hyperparameters are lightly tuned, and the average performance on the test set over 10 runs is reported along with the standard deviation. To enhance performance while mitigating the risk of overfitting, different sets of hyperparameters are used for tasks at different levels, considering the variation in sample sizes. For city-level electricity consumption prediction tasks, Ridge regression is adopted due to the limited number of data samples.
@@ 4.2 Performance Comparison
We benchmark MobCLIP against the strong existing approaches: 1) Self-supervised pertaining approaches: SatCLIP (with ViT-16) [19], GeoCLIP [38] and CSP [25]; and 2) analytical position encoding functions: Sphere2Vec (sphereC+) [26], Siren [33], and NeRF [29]. (The selection is based on methods’ superior performance reported in previous work [42]). Note that the location encoders used for comparison in this study take 2D coordinates as input, whereas MobCLIP operates on H3 grids as its spatial unit. To ensure consistency, we generate location embeddings for the centroids of the H3 grids for comparison with our grid embeddings.
@@
In Table 2, we compare the prediction performance of MobCLIP against baseline models across a range of tasks and organize the results based on the categories of downstream tasks. Clearly, MobCLIP achieves significantly better results across all human-related tasks, with an average improvement of 35.8% and 87.9% for social and economic tasks, respectively. In the two nature-related tasks (i.e., elevation and forest coverage prediction), SatCLIP performs comparably or slightly better, showing a minor average difference of 4.1 % in prediction performance. Note that both elevation and forest coverage data tend to exhibit gradual and continuous spatial changes, meaning that their variation is often highly correlated with the "absolute" location differences. Hence, it is reasonable that the location encoders with only latitude/longitude as input can capture the patterns well enough. However, the majority of tasks that hold value in practice are related to human behavior, which typically does not exhibit easily inferable location-dependent properties. In such tasks, MobCLIP performs significantly better, showcasing its ability to capture complex, high-level region correlations beyond simple geographic proximity. Meanwhile, the minor performance gap in the two nature-related tasks compared to top-performing methods suggests that our model also effectively captures direct region adjacency along with more nuanced features.
Table 2: Comparison of model performances on 11 downstream tasks spanning social, economic, and natural aspects. Previous best results are underlined, while best results are highlighted in bold. Model performances are measured by $R^{2}$, and the standard deviation over ten repeated trainings are marked in brackets below. Row $\Delta$ illustrates the percentage difference between Our Method’s performance and the best baseline.
Social Tasks Economic Tasks Natural Tasks
@@
256 0.52 0.63 0.68 0.74 0.39 0.33 0.66 -0.07 0.44 0.96 0.81 0.55
SatCLIP [19] (0.03) (0.01) (0.03) (0.03) (0.01) (0.03) (0.03) (0.11) (0.03) (0.00) (0.01)
512 0.41 0.66 0.66 0.69 0.32 0.24 0.65 0.11 0.32 0.92 0.73 0.52
GeoCLIP [38] (0.04) (0.01) (0.02) (0.03) (0.01) (0.03) (0.02) (0.01) (0.00) (0.00) (0.01)
256 0.55 0.65 0.62 0.68 0.39 0.29 0.62 0.20 0.46 0.95 0.77 0.55
CSP [25] (0.03) (0.02) (0.03) (0.03) (0.01) (0.03) (0.02) (0.03) (0.03) (0.00) (0.01)
96 0.40 0.56 0.62 0.61 0.34 0.26 0.57 -0.01 0.35 0.95 0.76 0.49
Sphere2Vec [26] (0.03) (0.02) (0.04) (0.03) (0.01) (0.03) (0.02) (0.01) (0.03) (0.00) (0.01)
96 0.40 0.63 0.61 0.61 0.33 0.24 0.56 0.03 0.33 0.95 0.76 0.50
NeRF [29] (0.03) (0.02) (0.03) (0.03) (0.01) (0.03) (0.02) (0.05) (0.02) (0.00) (0.01)
1,024 0.51 0.66 0.69 0.74 0.39 0.33 0.66 -0.14 0.44 0.96 0.80 0.54
Siren [33] (0.03) (0.02) (0.03) (0.03) (0.01) (0.03) (0.03) (0.14) (0.03) (0.00) (0.01)
@@
Additionally, the superiority of MobCLIP is consistent across varied spatial scales, achieving the best performance on tasks at city, county, and grid levels. Notably, in the city-level energy consumption task, MobCLIP demonstrates impressive results with a prediction accuracy of 0.72 in terms of $\text{R}^{2}$. In contrast, the baseline models exhibit $\text{R}^{2}$ values fluctuating around zero, failing to capture regional features at coarser resolutions.
@@ 4.3 Scaling Behavior in Geospatial Representation Learning
Echoing scaling behaviors observed in LLMs [16], we investigate the scaling of model performance in geospatial representation learning with pretraining data size and spatial coverage. Figure 2 illustrates model performances of three MobCLIP variants, respectively trained using data from three nested spatial coverages, namely Jiangsu, East China, and China. The spatial extents are shown in Figure 2(r). We conducted three separate experiments with different pretraining data sizes. While Jiangsu contains 3,904 H3 cells, East China contains 28,855 cells and China contains 195,574 cells. All three pretrained models are tested in Jiangsu only, using evaluation methods described in Section 4.1.2.
We observe that downstream model performance generally increases with increased pretraining data size for both social and economic tasks (Figure 2, Appendix A.2). The performance increase from Jiangsu to East China is especially large, while the increase from East China to China is relatively marginal. External training information can be useful for internal predictions. The marginal benefits of increasing training data size gradually decline. The scaling behavior in geospatial representation learning is consistent with findings in language and vision domains, where model performance improves with scale [13]. Results emphasize the value and necessity of geospatial representation learning at scale. Even for small-scale predictions, an increase in training data size and spatial coverage will introduce relevant information and improve downstream predictive performances.
@@ 4.4 Ablation and Sensitivity Study
Figure 3(a) shows the performance of multiple model variants on downstream tasks for ablation analysis. First, the CLIP-based multimodal alignment module enriches the model by incorporating complementary information from various modalities. The columns labeled w/o CLIP indicate that removing this module leads to an average performance drop of  5% on human-related tasks. More notably, performance on natural-related tasks drops sharply from 85% to 63%, indicating that while mobility data is highly informative, it primarily captures human-centric patterns. In contrast, multimodal inputs—such as remote sensing imagery—provide richer representations of natural features. These results show that integrating diverse modalities leads to a more comprehensive and general-purpose regional representation. Second, mobility forms the backbone of MobCLIP, and graph information is integrated at two distinct stages. We evaluate each stage by removing them individually. When the graph encoder is removed, we replace it with a simple MLP. As shown by w/o Mob Graph Encoder, the performance drop is most pronounced on human-related tasks, suggesting the GCN effectively captures human-centric patterns via relational structure. Replacing LINE initialization with random values (w/o LINE Initialization) has the smallest impact across task types. This suggests LINE is less effective than GCN in capturing informative structure. However, a consistent 1–2% drop across tasks indicates that LINE still provides complementary value.
@@
Considering the importance of the mobility graph encoder, we conduct a sensitivity analysis on both the node neighbours sampling strategy and the sampling ratio (Figure 3(b)). We compare our top-$k$ sampling approach with random sampling, which retains links randomly based on a percentage threshold instead of depending on link weights. The results suggest that varying the top-$k$ sampling ratio has limited effect on performance, as the difference between top 10% and top 30% sampling remains marginal. In contrast,random sampling tends to include nodes from the "tail" of the long-tail distribution, introducing noise that adversely affects model accuracy. As a result, random sampling generally underperforms top-$k$ sampling at equivalent ratios. However, as the random sampling ratio increases, the model gradually captures broader and more generalized patterns, while the relative impact of noise diminishes, leading to improved performance. These findings highlight that MobCLIP is a unified framework capable of balancing performance and computational efficiency under different graph configurations.
5 Conclusion

### v4

@@ 4.1.1 Multimodal Pretraining
We construct a store-level network based on consumer mobility from WeChat Pay, a nationwide electronic transaction platform in China operated by Tencent, with extensive offline payment coverage. Each offline store is mapped to an H3 grid cell by coordinates. If a user transacts at stores in different grids within the same week, we add an undirected, weighted edge between those cells, forming a weekly mobility graph. Aggregating 54 weeks across 50 million stores yields a large-scale graph at H3 resolution level 6, with approximately 200,000 nodes and 1.2 billion edges.
@@
We leverage 100M+ Points-of-Interest (POIs) from Tencent Maps, each with a name and category label spanning 445 categories (Tencent LBS Platform, 2024). POI names are typically short, referring to store names, company names, or types of public facilities. For imagery, we fetch Google Satellite imagery at approximately 10m resolution (zoom level 14) for each H3 grid. Demographic statistics are derived from WorldPop (WorldPop, 2018) by aggregating population count per H3 grid, stratified by 10-year age bands, gender, and work-home status in 2020.
@@
We split the dataset into 90% for training and 10% for validation to mitigate overfitting. The model is trained for 100 epochs with a batch size of 20,480, using the Adam optimizer with a learning rate of 0.001 and weight decay of 0.001 on a single NVIDIA A100 GPU for one hour. During training, we conduct a grid search over commonly used hyperparameters: learning rate $\in\{1e-4,2e-4,5e-4,1e-3,2e-3,5e-3\}$ and weight decay $\in\{1e-3,1e-2,1e-1\}$. The LINE embeddings for the nationwide mobility graph were computed on Tencent’s Angel distributed graph computing platform (Jiang et al., 2018), utilizing hundreds of computing cores. Graph nodes with more than a billion edges were embedded efficiently within a few hours.
4.1.2 Downstream Evaluation
@@
To assess our model’s generalizbility, we curated a set of 9 downstream tasks as shown in Table 1. These tasks were selected for their diversity and broad representativeness, spanning the two key domains of human‑centric scenarios: social and economic. Considering the influence of statistical unit delineation in geospatial analysis, the downstream tasks are designed to span four distinct spatial scales — point, grid, county/district, and city/prefecture — to examine resolution effects in region representation learning. The samples are evenly distributed across the study area, ensuring an evaluation free of geographic bias. Detailed descriptions and spatial distributions of the dataset samples are provided in Appendix A.2. MoRA is trained at the H3 grid level. For point-level tasks, individual point values within each grid are aggregated to align with the grid-level embeddings. For administrative-level tasks, we average embeddings over all grids within each unit and use the resulting mean for regression. This multi-scale evaluation allows us to assess the model’s generalizability across varying spatial scales.
@@
Abbreviation Description Data Source Spatial Scale Sample number Publicly Available
@@
POP Population density WorldPop (Bondarenko et al., 2020) Grid 194,232 Y
EDU Average education years Census Year Book (NBS, 2020) County/District 2,800 Y
ELD Elderly (65+) population ratio Census Year Book (NBS, 2020) County/District 2,800 Y
HSR Hukou separation ratio Census Year Book (NBS, 2020) County/District 2,800 Y
CRI Crime cases Zhang et al. (2025) Grid 39,932 Y
@@
NTL Nighttime light intensity MOSAIKS (Rolf et al., 2021) Point 23,333 Y
HOU Per capita housing area Census Year Book (NBS, 2020) County/District 2,800 Y
ENE Energy consumption City statistical yearbooks City/Prefecture 283 N
COS Offline consumption amount Tencent Grid 194,052 N
@@
We implement LightGBM(Ke et al., 2017) as the downstream prediction model. The hyperparameters are lightly tuned, and the average performance on the test set over 10 runs is reported along with the standard deviation. To enhance performance while mitigating the risk of overfitting, different sets of hyperparameters are used for tasks at different levels, considering the variation in sample sizes. For city-level electricity consumption prediction tasks, Ridge regression is adopted due to the limited number of data samples.
@@ 4.2 Performance Comparison
We benchmark MoRA against the strong existing approaches from two categories:
  1. 1.
    Approaches with publicly available, pre-trained embeddings, which we evaluate directly on our benchmark tasks and datasets. These can be further grouped into: 1) Self-supervised pretraining approaches: AlphaEarth (Brown et al., 2025), SatCLIP (with ViT-16) (Klemmer et al., 2025), GeoCLIP (Vivanco et al., 2023), and CSP (Mai et al., 2023a); and 2) analytical position-encoding functions: Sphere2Vec (sphereC+) (Mai et al., 2023b), Siren (Rußwurm et al., 2024), and NeRF (Mildenhall et al., 2020). (The selection is based on methods’ superior performance reported in previous work (Wu et al., 2024)). Note that MoRA operates on H3 grids as its spatial unit. To ensure consistency, for the location encoders used for comparison that take 2D coordinates as input, we generate location embeddings for the centroids of the H3 grids for comparison with our grid embeddings. For the pixel-level encoder, i.e., AlphaEarth, the mean of all pixel embeddings within each grid is computed as the baseline. We use grids covering the entirety of China for comparison.
  2. 2.
    Approaches that are not directly applicable, requiring further work to generate embeddings, either by training a published model on our dataset or by sourcing compatible input data to perform inference with their pre-trained weights. Since most existing models from this type were designed and tested at the city scale instead of a national one, we implement all models in Jiangsu Province for a fairer comparison (see Figure 2 for Jiangsu’s location). We select HREP (Zhou et al., 2023), ReCP (Li et al., 2024), and ReFound (Xiao et al., 2024) as representative works from each type as baselines. See Appendix A.3 for details of the baselines and our implementation for comparison.
@@
In Table 2, we compare MoRA’s predictive performance against pre-trained embeddings across a range of tasks and report results by downstream task category. Clearly, MoRA delivers superior performance across all tasks, averaging a 12.9% improvement overall, with 10.8% and 16.0% gains for the social and economic tasks, respectively. Note that the very slight underperformance on nighttime-light prediction (1.7%) likely stems from implicit correlations and co-variation between common Earth observation features and the target indicator. Given that human-centric tasks hold great value in practice and typically do not exhibit easily inferable location-dependent properties, an average $R^{2}$ of 0.78 across nine diverse tasks demonstrates that MoRA effectively captures complex, high-level regional correlations beyond simple geographic proximity, delivering strong practical value.
Table 2: Comparison of model performances on 9 downstream tasks spanning social and economic aspects. Previous best results are underlined, while best results are highlighted in bold. Model performances are measured by $R^{2}$, and the standard deviations over ten repeated trainings are marked in brackets below. Row $\Delta$ illustrates the percentage difference between Our Method’s performance and the best baseline.
Social Tasks Economic Tasks
@@
64 0.80 0.77 0.71 0.68 0.71 0.63 0.63 0.47 0.81
AlphaEarth (0.02) (0.01) (0.02) (0.03) (0.01) (0.02) (0.03) (0.02) (0.01)
256 0.52 0.63 0.68 0.74 0.39 0.33 0.66 -0.07 0.44
SatCLIP (0.03) (0.01) (0.03) (0.03) (0.01) (0.03) (0.03) (0.11) (0.03)
512 0.41 0.66 0.66 0.69 0.32 0.24 0.65 0.11 0.32
GeoCLIP (0.04) (0.01) (0.02) (0.03) (0.01) (0.03) (0.02) (0.01) (0.00)
256 0.55 0.65 0.62 0.68 0.39 0.29 0.62 0.20 0.46
CSP (0.03) (0.02) (0.03) (0.03) (0.01) (0.03) (0.02) (0.03) (0.03)
96 0.40 0.56 0.62 0.61 0.34 0.26 0.57 -0.01 0.35
Sphere2Vec (0.03) (0.02) (0.04) (0.03) (0.01) (0.03) (0.02) (0.01) (0.03)
96 0.40 0.63 0.61 0.61 0.33 0.24 0.56 0.03 0.33
NeRF (0.03) (0.02) (0.03) (0.03) (0.01) (0.03) (0.02) (0.05) (0.02)
1,024 0.51 0.66 0.69 0.74 0.39 0.33 0.66 -0.14 0.44
Siren (0.03) (0.02) (0.03) (0.03) (0.01) (0.03) (0.03) (0.14) (0.03)
@@
Additionally, the superiority of MoRA is consistent across varied spatial scales, achieving the best performance on tasks at city, county, and grid levels. Notably, in the city-level energy consumption task, MoRA demonstrates impressive results with a prediction accuracy of 0.72 in terms of $\text{R}^{2}$. In contrast, the baseline models exhibit $\text{R}^{2}$ values mostly fluctuating around zero, while our approach exceeds the second-best by 54.5%, revealing their limited capacity to capture regional features at coarser scales.
For the second category of approaches, Table 3 shows that MoRA, with a compact 128-dimensional representation, consistently outperforms both baselines, featuring an average gain of 20.9% over the second-best. This underscores MoRA’s methodological advantages as a generalizable framework.
Table 3: Performance comparison between MoRA and methodological baselines in Jiangsu Province. The $\Delta$ row reports the percentage improvement of our method over the strongest baseline.
[…]
POP EDU ELD HSR CRI NTL HOU ENE COS
HREP 128 0.57 0.55 0.24 0.49 0.68 0.34 0.42 0.38 0.73
ReCP 96 0.34 0.78 0.34 0.55 0.73 0.45 0.33 0.26 0.64
ReFound 768 0.41 0.71 0.45 0.51 0.73 0.47 0.35 0.80 0.68
MoRA 128 0.68 0.75 0.55 0.81 0.74 0.55 0.53 0.73 0.85
[…]
Inference.
For practical deployment, we develop and open-source a distilled version of MoRA that captures core geospatial knowledge and enables direct, coordinate-only inference. Specifically, we train a multi-layer perceptron (MLP) to map continuous geographic coordinates to high-dimensional embeddings, supervised by MoRA’s grid-level outputs. This allows practitioners to generate high-quality, privacy-preserving embeddings at arbitrary locations using only geographic coordinates, completely eliminating the need for external datasets and H3-based lookups at inference time. Detailed implementation of the distillation approach is provided in Appendix A.1.
@@ 4.3 Scaling Behavior in Geospatial Representation Learning
Echoing scaling behaviors observed in LLMs (Kaplan et al., 2020), we examine how model performance in geospatial representation learning scales with pretraining data size and spatial coverage. Figure 2 shows model performances of three MoRA variants trained on three nested spatial coverages, namely Jiangsu, East China, and China. The spatial extents are shown in Figure 2(r). We conducted three separate experiments with different pretraining data sizes where Jiangsu contains 3,904 H3 cells, East China contains 28,855 cells and China contains 195,574 cells. All three pretrained models are tested in Jiangsu only, using evaluation methods described in Section 4.1.2.
We observe that downstream performance generally increases with larger pretraining data for both social and economic tasks (Figure 2, Appendix A.4). The performance gain from Jiangsu to East China is especially large, while the increase from East China to China is relatively marginal. External training information can be useful for internal predictions. The marginal benefits of increasing training data size gradually decline. The scaling behavior in geospatial representation learning is consistent with findings in language and vision domains, where model performance improves with scale (Henighan et al., 2020). Results emphasize the value and necessity of geospatial representation learning at scale. Even for small-scale predictions, an increase in training data size and spatial coverage will introduce relevant information and improve downstream predictive performances.
@@ 4.4 Ablation and Sensitivity Study
Ablation study on anchoring modality.
We frame mobility as the backbone for aligning other modalities and justify this design theoretically. Here, we empirically evaluate its unique advantages over other anchoring choices. As shown in Figure 3(a), mobility-as-backbone significantly outperforms alternative anchor modalities across both task types, achieving an average of 16.8% gain over the second-best (the demo-anchored structure), validating its superiority.
@@
Figure 3(b) shows the performance of multiple model variants on downstream tasks. The columns labeled w/o CLIP indicate that removing the CLIP-based alignment module and leaving only the mobility modality leads to an average performance drop of 6.5%. Second, mobility forms the backbone of MoRA, with graph information injected at two stages. We assess each by ablating them separately. Removing the graph encoder and substituting a simple MLP (w/o Mob Graph Encoder) causes the largest drop, indicating the GCN captures non-local patterns via relational structure. Replacing LINE initialization with random values (w/o LINE Initialization) has a smaller impact, suggesting that the graph structure is more informative than pre-encoded embeddings. However, a consistent 1–2% decline shows that LINE still provides complementary value. We also perform an ablation on each modality to evaluate their contributions (Appendix A.5). We further replace the mobility graph with a simple neighborhood graph in which edges connect only directly adjacent cells (MoRA(neighborhood graph)). This leads to a 12.2% drop in average performance, highlighting the mobility graph’s ability to capture long-range spatial correlations, i.e., ”functional adjacency”, that are absent in graphs based solely on direct neighborhood adjacency.
@@
Given the mobility graph encoder’s importance, we analyze sensitivity to neighbor sampling strategy and ratio (Figure 3(c)). We compare our top-$k$ sampling approach with random sampling, which retains links randomly based on a percentage threshold instead of depending on link weights. The results suggest that top-$k$ sampling shows minimal performance variation across ratios (top 10% vs. top 30%). Random sampling, by tending to include nodes from the ”long-tail”, introduces noise and generally underperforms top-k at the same ratios. As a result, random sampling generally underperforms top-$k$ sampling at equivalent ratios. Yet as the random ratio increases, more generalized patterns offset the noise, leading to improved performance. Overall, MoRA balances accuracy and efficiency across graph configurations.
5 Conclusion and Discussion